OpenStudy (anonymous):

Find two numbers whose product is -40, and whose sum is 3

OpenStudy (amistre64):

well; (2+n) + (1-n) = 3 (2+n)(1-n) = -40 ... solve for n to determine the numbers ....

OpenStudy (anonymous):

6?

OpenStudy (amistre64):

6 feels good ... 8-5 = 3, and 8(-5) = -40

OpenStudy (anonymous):

And -7 for the first part?

OpenStudy (amistre64):

(2+n) + (1-n) = 3 6 6 8 + (-5) = 3, is good (2+n)(1-n) = -40 6 6 8 (-5) = -40 this is a method that some guy in greece used ... i think his name was diophantus